33. Suppose that King Kanuta, whom you met in your workbook, demands that each of his subjects give him I coconut for every coconut that they consume. The king puts all of the coconuts that he collects in a large pile and burns them. The supply of coconuts is given by Sup)100p., where p, is the price received by suppliers. The demand for coconuts by the kings subjects is given by Dipa) -1,500 100pa where pa is the price paid by consumers. In equilibrium, the price received by suppliers will be Pa
Answer #1

To find Equilibrium price we need to equate demand and supply equation as,

Qd=Qs

1500-100p=100P

1500=200P

P = 1500/200= $7.5